. A factory produces very small electric light bulbs. It is known from experience that 6% of the bulbs produced are defective, i.e. they do not work.
Another factory uses these bulbs in strings of lights used to decorate Christmas trees, where each string contains 50 light bulbs.
For a string of lights, calculate,
(a) the mean number of bulbs which are defective







(b) the variance of the number of defective bulbs





(c) the standard deviation of the number of defective bulbs






(d) The company that produces the strings of lights packages each string in a box. The company also wishes to include a small packet of spare bulbs in the box. Using your answers from (a), (b) and (c) above, how many spare bulbs should they include with each string?


Answer by stanbon(75887)

A factory produces very small electric light bulbs. It is known from experience that 6% of the bulbs produced are defective, i.e. they do not work.
Another factory uses these bulbs in strings of lights used to decorate Christmas trees, where each string contains 50 light bulbs.
For a string of lights, calculate,
(a) the mean number of bulbs which are defective
np = 50*0.06 = 3
------
(b) the variance of the number of defective bulbs
npq = 3*0.4 = 1.2
------
(c) the standard deviation of the number of defective bulbs
sqrt(npq) = sqrt(1.2) = 1.095..
------
(d) The company that produces the strings of lights packages each string in a box. The company also wishes to include a small packet of spare bulbs in the box. Using your answers from (a), (b) and (c) above, how many spare bulbs should they include with each string?
Guess: 53
-----
Cheers,
Stan H.